OpenStudy (anonymous):

Use the quadratic formula to solve equation 2x to the power of 2 -3x+10=0

OpenStudy (mathteacher1729):

This is the question? \[\huge 2x^{2-3x+10}=0\] Or is it \[\huge 2x^{x^2-3x+10}=0\]

OpenStudy (anonymous):

2x^2-3x+10=0

OpenStudy (anonymous):